OpenStudy (anonymous):

find the equation of the circle that passes through the origin and has intercepts equal to 1 and 2 on the x and y axes, respectively.

OpenStudy (anonymous):

two points (1,0) ,(0,2)

OpenStudy (anonymous):

WRITE d eqn of circle ..den put 0,0 den 1,0 and den 0,2 ..simple

OpenStudy (anonymous):

(x-a)^2 +(y-b)^2=r^2 three equations (-a)^2 +(-b)^2=r^2 (1-a)^2 +(-b)^2=r^2 (-a)^2 +(2-b)^2=r^2

OpenStudy (anonymous):

Don't say its simple, that's degrading, and don't write den instead of then.

OpenStudy (anonymous):

I got a=1/2 , b=1 , r^2= 5/4

OpenStudy (anonymous):

is the answer \[(x-\frac{1}{2})^2+(y-1)^2=\frac{25}{4}\]?

OpenStudy (anonymous):

oops yes you are right imran sorry

OpenStudy (anonymous):

I got \[\left(x-\frac{1}{2}\right)^2+(y-1)^2=\frac{5}{4}\]